OpenStudy (anonymous):

Players A and B each have a well shuffled standard pack of cards, with no jokers. The players deal their cards one at a time, from the top of the deck, checking for an exact match. Player A wins if, once the packs are fully dealt, no matches are found. Player B wins if at least one match occurs. What is the probability that player A wins?

OpenStudy (anonymous):

you gotta be kidding. this is a real pain

OpenStudy (anonymous):

I am not kidding.

OpenStudy (anonymous):

No googling allowed!

OpenStudy (anonymous):

oh it is a quiz!

OpenStudy (dumbcow):

haha how are you gonna enforce that...that tells me this question is out there somewhere huh? :)

OpenStudy (anonymous):

OUI!YES!

OpenStudy (dumbcow):

can i write a computer program to simulate the game, then determine the probability

OpenStudy (anonymous):

Well you can do that.

OpenStudy (anonymous):

But you can also use your brain instead of a CPU! :)

OpenStudy (anonymous):

yes the answer is out there, as you might imagine it is a standard problem, and the solution is cute

OpenStudy (dumbcow):

is the probability that on any given turn they have a match 1/52 ?

OpenStudy (dumbcow):

well my guess is \[(\frac{51}{52})^{52} = 0.364\]

OpenStudy (anonymous):

Use the inclusion - exclusion principle.You'll get a surprising result!

OpenStudy (anonymous):

For any reasonably large number of cards, say, 10 or more, the probability that A wins is almost independent of the number of cards in the decks!

OpenStudy (anonymous):

The probability that A wins is d(52) / 52! = 1 − 1/1! + 1/2! − 1/3! + ... + 1/52!

OpenStudy (anonymous):

or,1/e

OpenStudy (dumbcow):

ah ok, well its time to get off OS oh solution thanks...hey i was close

OpenStudy (anonymous):

Yeah!You were!

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=1%2Fe&t=crmtb01

OpenStudy (dumbcow):

just wanted to add something i noticed \[\lim_{n \rightarrow \infty}(\frac{n-1}{n})^{n} = \frac{1}{e}\] so 1/e is a good approximation for large n but the exact solution for deck of 52 would be (51/52)^52 no?

OpenStudy (anonymous):

@dumbcow i think your answer, assuming it is right and it certainly looks right, is much simpler than the answer given in the nicks puzzles with all the "inclusion/exclusion" summing over stuff business i am wondering why, because your solution sure looks correct and straightforward, or am i missing something?

OpenStudy (dumbcow):

i don't know much about inclusion/exclusion honestly i wasn't really trying, but when he said solution was 1/e it made me pretty sure my answer is the correct one for a deck of 52 cards logic: chance of a specific card is drawn is 1/52 for each player probability they match = 1/52 * 1/52 = 1/52^2 there are 52 different cards so probability any given card matches = 52/52^2 = 1/52 probability they don't match = 51/52 for 52 trials --> (51/52)^52